OpenStudy (anonymous):

A lamp costs $87.75. If the sales tax is 8%, how much tax would be paid on this lamp?

OpenStudy (anonymous):

help me

hartnn (hartnn):

8% of 87.75 = (8/100)* 87.75 = ...

OpenStudy (anonymous):

i did not understand

hartnn (hartnn):

x% = x/100 so, 8% = 8/100 got this ?

OpenStudy (anonymous):

yah

hartnn (hartnn):

so, 8% of 87.75 = (8/100)* 87.75 calculate.

OpenStudy (anonymous):

@hartnn is rite

OpenStudy (anonymous):

yah you found me ilovehomework

OpenStudy (anonymous):

Grab your calculator. Multiply $87.75 by 0.08. So, it would be $7.02

OpenStudy (anonymous):

so is that the answer

hartnn (hartnn):

yes, it is.

OpenStudy (anonymous):

thank you